# SAMM: Sharded Automated Market Makers

### v1

@@ Appendix G Equilibrium Proof Details
G.7 Proof of Theorem 7.2.4
  • Theorem 7.2.4 (restated). In $\Gamma_{n}(\textit{tf}_{SAMM})$, assume that the pool state in step $k$ is $\textbf{R}({k})$, $i^{\ast}=i_{\min}(\textbf{R}({k}))$ is the index of the pool with the smallest amount of deposited token A in $\textbf{R}({k})$. Then the trader strategy is to trade in the smallest pool in the last step if it is the smallest one, or randomly select one of the smallest pools, namely,
    $\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}1,&\text{if}R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\text{ and }\\ &a^{\textit{BA}}=a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}})\\ \frac{1}{n_{\min}(\textbf{R})},&\text{if }R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\text{ and }\\ &a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases}$ (124)
@@
We first prove that $\hat{l}^{A}_{i^{\ast}}>0$ by contradiction. If $\hat{l}^{A}_{i^{\ast}}=0$, then $R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}=R^{A}_{i^{\ast}}$. Then for any input amount $\hat{l}^{A}_{j}>0$, we have $R^{A}_{j}+\hat{l}^{A}_{j}>R^{A}_{j}$. From the definition of $i^{\ast}=i_{\min}(\textbf{R})$ (Equation 7.2.4), we have $R^{A}_{i^{\ast}}\leq R^{A}_{j}$. Therefore, the amount of deposited token A of $\textit{pool}_{i^{\ast}}$ in $\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ is strictly smaller than $\textit{pool}_{j}$
@@
This contradicts the definition of $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ (Definition 7.2.1) since $\hat{l}^{A}_{j}>0$. Therefore, $\hat{l}^{A}_{i^{\ast}}>0$. Then also from that definition, the amount of deposited token A of $\textit{pool}_{i^{\ast}}$ after the fillup action is smallest among all pools:
@@
From the definition of $U_{lp}(\textbf{R},a_{lp},a^{\textit{BA}})$ (Equation 6.1.4), we have
@@
From Lemma 7.2.1, the minimal amount of deposited token A in $\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ is no less than that in $\textbf{R}+a_{lp}$ for any $a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})$:
@@
From Lemma 7.2.1, the trading fee of trading in a larger pool is larger:
$\textit{tf}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}),b^{\textit{BA}})>\textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\,\,.$ (134)
From the definition of $i^{\ast}=i_{\min}(\textbf{R})$ (Equation 7.2.4), we have $R^{A}_{i^{\ast}}\leq R^{A}_{i}$. Combining with Equation 133, we have
@@
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})$
@@
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})$
@@ Appendix H Evaluation of Increasing Parallel Component
We posit that enhancing the performance of SAMM necessitates mitigating the serial bottlenecks within the platform. While modifications to the core architecture of Sui are beyond the scope of this study, we enhance the parallelizable aspects of SAMM by introducing superfluous operations into each trade. This methodology follows the experimental setup described in Section 8.
Figure 9 presents the maximum throughput results for standard SAMM transactions as previously discussed in Section 8, alongside results from transactions with added operations with three repetitions. Although the inclusion of additional operations reduces overall performance due to increased overhead, it significantly enhances the parallel component, with $P=0.9$ ($R^{2}=0.968$). Remarkably, the throughput with $32$ shards is ten times that of a single shard, a substantial improvement over the ratios observed in Section 8.
@@ Appendix I Parameter Selection in Simulation
We select different values of $c$ for SAMM, namely $0.003$, $0.005$ and $0.01$. We set $r_{\max}=5\times r_{\min}$ We choose $r_{\max},r_{\min}$ and $\beta_{1}$ (see Section 5.3) to minimize the maximal trading fee ratio. This is to limit the maximal cost for traders. Hence, we set $r_{\max}=5\times r_{\min}$ to minimize them at the same time. Together with the restrictions of satisfying $c$-smaller-better and $c$-larger-better (Corollary 5.3), the optimization problem is

### v2

@@ G-B3 Convergence to Balanced Pools
G-B4 Specific SPNE under deviation
The above theorem does not prove the existence of an SPNE. We settle this by finding a specific SPNE in $\Gamma_{n}(\textit{tf}_{SAMM})$. In this SPNE, if there are multiple smallest pools, the trader uses the smallest pool in the last step. If there is more than one smallest pool in the last step, the trader selects the one with the smallest index. Denote by $i_{\min}(\textbf{R})$ the index of the pool with the smallest amount of deposited token A in R. When $i^{\ast}=i_{\min}(\textbf{R})$, we have
$\forall j,R^{A}_{i^{\ast}}\leq R^{A}_{j}$
$\forall j,R^{A}_{i^{\ast}}=R^{A}_{j}\Rightarrow i^{\ast}\leq j\,\,.$ (123)
If the pool with index $i_{\min}(\textbf{R})$ is the only pool, the BA trader would only trade in that pool. Then the liquidity provider only taking the fill-up action is the best response strategy:
{thm}
[] In $\Gamma_{n}(\textit{tf}_{SAMM})$, assume that the pool state in step $k$ is $\textbf{R}({k})$, $i^{\ast}=i_{\min}(\textbf{R}({k}))$ is the index of the pool with the smallest amount of deposited token A in $\textbf{R}({k})$. Then the trader strategy to trade in the smallest pool in the last step if it is the smallest one, or randomly select one of the smallest pools, namely,
$\tau^{BA}(\textbf{R},b^{\textit{BA}},a^{\textit{BA}})=\begin{cases}1,&\text{if}R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\\ &\text{and }a^{\textit{BA}}=a_{i^{\ast}}^{\textit{BA}}(b^{\textit{BA}})\\ \frac{1}{n_{\min}(\textbf{R})},&\text{if }R_{i^{\ast}}=\rho^{A}(\textbf{R}({k}),l^{A},l^{B})\\ &\text{and }a^{\textit{BA}}\in\mathcal{A}_{\mathbf{1},\min}(b^{\textit{BA}},\textbf{R})\\ 0,&\text{Otherwise.}\end{cases}$ (124)
@@ G-B4 Specific SPNE under deviation
We first prove that $\hat{l}^{A}_{i^{\ast}}>0$ by contradiction. If $\hat{l}^{A}_{i^{\ast}}=0$, then $R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}=R^{A}_{i^{\ast}}$. Then for any input amount $\hat{l}^{A}_{j}>0$, we have $R^{A}_{j}+\hat{l}^{A}_{j}>R^{A}_{j}$. From the definition of $i^{\ast}=i_{\min}(\textbf{R})$ (Equation G-B4), we have $R^{A}_{i^{\ast}}\leq R^{A}_{j}$. Therefore, the amount of deposited token A of $\textit{pool}_{i^{\ast}}$ in $\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ is strictly smaller than $\textit{pool}_{j}$
@@
This contradicts the definition of $a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ (Definition G-B1) since $\hat{l}^{A}_{j}>0$. Therefore, $\hat{l}^{A}_{i^{\ast}}>0$. Then also from that definition, the amount of deposited token A of $\textit{pool}_{i^{\ast}}$ after the fillup action is smallest among all pools:
@@
From the definition of $U_{lp}(\textbf{R},a_{lp},a^{\textit{BA}})$ (Equation VI-A4), we have
@@
From Lemma G-B1, the minimal amount of deposited token A in $\textbf{R}+a_{lp}^{\textit{fill}}(\textbf{R},l^{A},l^{B})$ is no less than that in $\textbf{R}+a_{lp}$ for any $a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})$:
@@
From Lemma G-B1, the trading fee of trading in a larger pool is larger:
$\textit{tf}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}},\frac{p^{A}}{p^{B}}(R^{A}_{i^{\ast}}+\hat{l}^{A}_{i^{\ast}}),b^{\textit{BA}})>\\ \textit{tf}(R^{A}_{i}+l^{A}_{i},\frac{p^{A}}{p^{B}}(R^{A}_{i}+l^{A}_{i}),b^{\textit{BA}})\,\,.$ (134)
From the definition of $i^{\ast}=i_{\min}(\textbf{R})$ (Equation G-B4), we have $R^{A}_{i^{\ast}}\leq R^{A}_{i}$. Combining with Equation 133, we have
@@
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\left(\begin{subarray}{c}\tau_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\\ \times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\end{subarray}\right)$
@@
$=$ $\sum_{a_{lp}\in\mathcal{A}_{lp}(l^{A},l^{B})}\left(\begin{subarray}{c}\pi_{lp}(\textbf{R},l^{A},l^{B},a_{lp})\\ \times U_{lp}(\textbf{R},a_{lp},\tau^{BA},\tau^{AB},l^{A},l^{B})\end{subarray}\right)$
@@
In summary, our equilibrium analysis indicates that the system can achieve a stable state where the perfect parallelism strategy is the dominant strategy for traders. Moreover, the system demonstrates robustness against deviations, including those caused by attacks.
@@ Appendix H Evaluation of Increasing Parallel Component
We posit that enhancing the performance of SAMM necessitates mitigating the serial bottlenecks within the platform. While modifications to the core architecture of Sui are beyond the scope of this study, we enhance the parallelizable aspects of SAMM by introducing superfluous operations into each trade. This methodology follows the experimental setup described in Section VII.
Figure 11 presents the maximum throughput results for standard SAMM transactions as previously discussed in Section 8, alongside results from transactions with added operations with three repetitions. Although the inclusion of additional operations reduces overall performance due to increased overhead, it significantly enhances the parallel component, with $P=0.9$ ($R^{2}=0.968$). Remarkably, the throughput with $32$ shards is ten times that of a single shard, a substantial improvement over the ratios observed in Section VII.
@@ Appendix I Parameter Selection in Simulation
We select different values of $c$ for SAMM, namely $0.003$, $0.005$ and $0.01$. We set $r_{\max}=5\times r_{\min}$ We choose $r_{\max},r_{\min}$ and $\beta_{1}$ (see Section V-C) to minimize the maximal trading fee ratio. This is to limit the maximal cost for traders. Hence, we set $r_{\max}=5\times r_{\min}$ to minimize them at the same time. Together with the restrictions of satisfying $c$-smaller-better and $c$-larger-better (Corollary V-C), the optimization problem is